OpenStudy (chihiroasleaf):

Given a series \( 4,x,y,z\) The first three terms are geometric series and the last three terms are arithmetic series. Find the value of \( x+y\)

OpenStudy (anonymous):

did u provide the value of contstants

OpenStudy (chihiroasleaf):

what do you mean?

Parth (parthkohli):

It's not a unique one.

Parth (parthkohli):

You can just add the \(z\) according to the values of \(x\) and \(y\).

OpenStudy (mathmath333):

Using the concept of arithmatic and geometric mean u get. \(\large \color{black}{\begin{align} x^2& =4y\hspace{.33em}\\~\\ 2y& =x+z\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (perl):

since the first 3 terms are geometric $$ \Large { \frac{ x} {4 }= \frac{y}{x} } $$ since the last three terms are arithmetic $$ \Large { y-x = z - y } $$

OpenStudy (chihiroasleaf):

I got stuck after that step :(

OpenStudy (perl):

$$ \Large { x^2 = 4y \\ 2y = z + x \\ \therefore \\ x^2 = 2(2y) \\ x^2 = 2( z + x)\\ x^2 -2x - 2z = 0 } $$ You can use quadratic formula to solve for x in terms of z

OpenStudy (chihiroasleaf):

I'll try...

OpenStudy (perl):

I assume the directions is to find x + y in terms of z

OpenStudy (anonymous):

i remember this one!

OpenStudy (mathmath333):

or try to put \(y=1,2,3\cdots \) there should be a pattern .

OpenStudy (chihiroasleaf):

I get \(x=1+\sqrt{1+2z}\) or \(x=1-\sqrt{1+2z}\) so, \(y=\frac{1+z+\sqrt{1+2z}}{2}\) or \(y=\frac{1-z+\sqrt{1+2z}}{2}\) so, \(x+y\) will be in term of \(z\), but this is multiple choice question, the options are -1 ; 0; 3; 7

OpenStudy (perl):

if its multiple choice its easier to plug in values sorry i didnt realize

OpenStudy (mathmath333):

u should clearly state if it was multiple choice or not otherwise it seems ambigous.

OpenStudy (perl):

did you get the answer @chihiroasleaf

OpenStudy (chihiroasleaf):

not yet :(

OpenStudy (perl):

ok so you have an expression for x + y , in terms of z. the solution is not unique, you can try plugging in values, the easiest value to plug in is z = 0

OpenStudy (chihiroasleaf):

if, z= 0, we get x+y = 3, and the sequence will be 4, 2, 1, 0 right?

OpenStudy (perl):

$$ \Large { x^2 = 4y \\ 2y = z + x \\ \therefore \\ x^2 = 2(2y) \\ x^2 = 2( z + x)\\ x^2 -2x - 2z = 0 \\ x = 1 + \sqrt{ 1+ 2z} , ~~1 - \sqrt{ 1 + 2z } \\ \therefore \\ y = \frac{z + x}{2} \\ \\ \therefore \\ y= \frac{z+1+\sqrt{1+2z}}{2} , ~ \frac{z+1-\sqrt{1+2z}}{2} \\ \therefore \\x + y\\ = (1 + \sqrt{ 1+ 2z}) + \frac{z+1+\sqrt{1+2z}}{2} \\ = \frac{2}{2}\cdot (1 + \sqrt{ 1+ 2z}) + \frac{z+1+\sqrt{1+2z}}{2} \\ \\ = \frac{2 + 2\sqrt{1 + 2z} + z + 1 + \sqrt{ 1 + 2z} }{2} \\ x+y = \frac{ 3\sqrt{1 + 2z} + z + 3 }{2} \\ \text { if z= 0 } \\ x+y = \frac{3 + 3} {2} = 3 } $$

OpenStudy (perl):

yes thats correct

OpenStudy (perl):

also z = 0 is a logical choice in order to avoid getting an irrational square root number

OpenStudy (perl):

z = 4 also produces a whole number output

OpenStudy (chihiroasleaf):

thanks a bunch @perl